# version 1.0

### v6

@@ 1.7 Bibliographic Remarks
Control theory is a mature and active discipline with both deep mathematical foundations and a rich history of applications. In terms of application, self-regulating feedback mechanisms that limit the flow of water, much like the modern float valve, were reportedly used in ancient Greece over two thousand years ago. The first mathematical analysis of such systems appeared in a paper by James Clerk Maxwell on the topic of governors [Max68], shortly after Maxwell had published his treatise on the unification of electricity and magnetism. For an enlightening historical account, see [FCZI03].
@@
Online nonstochastic control theory and the Gradient Perturbation Control algorithm were proposed in [ABH+19]. A flurry of novel methods and regret bounds for online nonstochastic control have appeared in recent years, many of which are surveyed in [HS21]. Notable among them are logarithmic regret algorithms [AHS19, FS20, Sim20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CH21], and nonstochastic control with partial observability [SSH20].
@@
In this chapter, we return to the general problem of control from Definition 1.2, that is, choosing a sequence of controls to minimize a sequence of cost functions. Before moving on to algorithms for control, it is important to reason about what kind of solution is even possible.
For this reason, we consider several natural examples of dynamical systems, and reason about what constitutes desirable behavior. This will help us understand the kinds of observation models and control objectives that are feasible and natural in real applications.
@@ 2.1.1 Medical Ventilator
In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m^{3}/s}$) into the lungs of a patient. This inflow approximately determines the volume of the lungs $v_{t}$ via a simple linear relation,
@@
where $\Delta$ is the time elapsed between successive actions of the discrete-time control algorithm. Being difficult to directly measure or estimate, the quantity $v_{t}$ is a hidden or latent variable from the controller’s perspective. Instead, the controller observes the air pressure $p_{t}$. We give a simplified physical model that relates the pressure and the volume. This model, given in the following, can be derived by first principles from stress-strain relations, treating the left and right lungs connected by the trachea as situated in a two-balloon experiment:
$p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3}$
The objective of the control algorithm is to follow an ideal pressure trajectory $p^{*}_{t}$ prescribed by the physician as closely as possible, say, as measured by squared loss $c_{t}(p_{t},u_{t})=|p_{t}-p^{*}_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort. See Figure 2.1 for a schematic representation of airflow when a lung is connected to a mechanical ventilator.
@@ 2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one-dimensional trajectory (i.e. the real number line). The state $\mathbf{x}_{t}$ is two dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity. The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step. The dynamics equation can be written as
@@
The position of the object is linearly related to its velocity via the parameter $\Delta$, which measures the time resolution the controller operates in. Optionally, the state is also affected by external perturbations $\mathbf{w}_{t}$ outside the influence of the controller.
The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{*}\in\mathbb{R}$ as quickly as possible, while also not expending too much energy. The cost function is naturally expressed as
$c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{*}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
@@
Figure 2.2: Double integrator illustration showing state coordinate $x_{1}(t)$ (position), state coordinate $x_{2}(t)$ (velocity), mass $m$, and control input $u(t)$.
@@ 2.1.3 Pendulum Swing-Up
The swing-up pendulum is one of the simplest examples of rotational kinematics. It corresponds to a point object of mass $m$ mounted at the end of a massless rod of length $l$. The other end of the rod is fixed to a pivot. Being vertically aligned, the system is subject to gravity with gravitational acceleration $g$. The controller is responsible for applying a torque $u_{t}\in\mathbb{R}$ at each time step. Given the vertical angle $\theta_{t}$ of the rod, the physics of this system may be summarized as
$\theta_{t+1}=\theta_{t}+\Delta\dot{\theta}_{t},\qquad\dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\frac{u_{t}-mg\ \sin\theta_{t}}{ml^{2}},$
where $\Delta$ is the time resolution of the controller. Note that this dynamical equation is nonlinear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the term $\sin\theta$ in the evolution of $\dot{\theta_{t}}$.
PivotmmAngle $\theta_{t}$$m\mathbf{g}$$\mathbf{u}_{t}$$l$
Figure 2.3: Pendulum swing-up illustration with marked angle $\theta_{t}$, gravitational force mgmg, control input torque $u_{t}$, mass $m$, and rod length $l$.
The objective here is to get the rod to rest in a vertical upright position at an angle of $\pi$, using as little torque as possible throughout the process. This may be expressed through the cost function
@@
2.1.4 Aircraft Dynamics
A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers. We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10]. The two dimensional control input $\mathbf{u}_{t}$ corresponds to the elevation angle and the thrust. The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
The following dynamics equation has matrices whose numerical values were computed using the dynamics equation for an actual aircraft. These numerical values can be determined by the mass, length, and other physical properties of the aircraft:
$\mathbf{x}_{t+1}=$ $\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\\ 0.020&-0.101&-0.429&0\\ 0&0&1&0\end{bmatrix}\mathbf{x}_{t}+\begin{bmatrix}0.01&1\\ -0.18&-0.04\\ -1.16&0.598\\ 0&0\end{bmatrix}\mathbf{u}_{t}$
$+\begin{bmatrix}-0.003&0.039&0&-0.322\\ -0.065&-0.319&7.74&0\end{bmatrix}\mathbf{w}_{t}$
Figure 2.4: Longitudinal control in an aircraft. Graphic by User_A1, licensed under $C$-BY-SA 4.0; it is an SVG derivative of a public domain work by G J Vigurs.
The quantities of interest for assessing the performance of the controller are the aircraft speed, climb rate — each of which is a linear transformation of the state — and applied thrust. The corresponding cost is
$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\begin{bmatrix}1&0&0&0\\ 0&-1&0&7.74\end{bmatrix}\mathbf{x}_{t}-\mathbf{x}_{t}^{*}\right\|^{2}+\left\|\begin{bmatrix}0&1\end{bmatrix}\mathbf{u}_{t}\right\|^{2}.$
@@ 2.1.5 Epidemiological Models
The COVID-19 pandemic marked an increase in public interest in policy discussions on the effectiveness of certain interventions as a means of controlling infection spread. Epidemiological models predict the spread of an infectious disease via structural models and differential equations. We present a discrete-time analogue of the most basic of such models, the SIR model. Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the population sizes of those susceptible, infected, and recovered, respectively. Let $u_{t}$ be the number of vaccines administered at time $t$. The dynamics of the basic SIR model are given by
$S_{t+1}$ $=S_{t}-\beta S_{t}I_{t}-\alpha u_{t}$
$I_{t+1}$ $=I_{t}+\beta S_{t}I_{t}-\gamma I_{t}$
$R_{t+1}$ $=R_{t}+\gamma I_{t}.$
Here, the coefficient $\beta$ measures the level of communicability of the infection, $\gamma$ measures the rate at which an infected person recovers, and $\alpha$ represents the efficacy of the vaccine in preventing infections.
Figure 2.5: Trajectory of a continuous-time SIR model with parameters $\alpha=0,\beta=0.3,\gamma=0.1$.
@@ 2.2 Solution Concepts for Dynamical Systems
Before considering the control, or manipulation, of a dynamical system, we must first reason about what the goal should even be. In this section, we discuss various solution concepts for a dynamical system, including equilibrium, stability, stabilizability, and controllability.
Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, the equivalent deterministic dynamical system, defined through the dynamics $\mathbf{x}_{t+1}=f^{\prime}(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$, models the evolution of the dynamical system in the absence of perturbations. In addition, the autonomous dynamical system $\mathbf{x}_{t+1}=f^{\prime\prime}(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$ describes the state transitions of the dynamical system in the absence of both control inputs and perturbations.
@@ 2.2.1 Equilibrium of Dynamical Systems
An example of an equilibrium is a ball sitting at rest at the bottom of a pit. If the ball were to be slightly moved, it would return to the bottom of the pit. Thus, this equilibrium is stable and attractive.
If the ball is at rest on the top of a mountain, pushing the ball in any direction would cause it to fall. This is an equilibrium that is unstable. A formal definition is as follows.
@@
A point $\mathbf{x}$ is an equilibrium point of an autonomous dynamical system $f$ if and only if $\mathbf{x}=f(\mathbf{x})$. An equilibrium point $\mathbf{x}$ is stable and attractive if and only if $\mathbf{x}=\lim_{t\rightarrow\infty}f^{t}(\mathbf{x}_{0})$ for every initial state $\mathbf{x}_{0}$ in some neighborhood of $\mathbf{x}$. If the neighborhood includes the entire set of states, we say that the equilibrium is globally attractive.
@@
Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all A{A}. If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vector converges to the null vector. Thus, $\mathbf{0}$ is a globally attractive and stable equilibrium.
In general, determining whether a system has an equilibrium is an intractable problem, as we show in this chapter.
@@ 2.2.2 Stabilizability and Controllability
Two fundamental properties of dynamical systems are stabilizability and controllability, as defined next.
@@
A dynamical system is stabilizable if and only if, for every initial state, there exists a sequence of controls that drives the deterministic-equivalent system to the zero state. That is, $\lim_{t\rightarrow\infty}{\mathbf{x}_{t}}=\mathbf{0}$.
@@
A dynamical system is controllable if and only if, for every initial and target state, there exists a sequence of controls that drives the equivalent deterministic system to the target state.
Note that controllability clearly implies stabilizability, as the zero state is one possible target state.
In the next section, we discuss computational difficulties that arise in determining if a system is stabilizable and/or controllable.

### v8

# version 1.2 $\beta$

@@ 1.6.1 Online Nonstochastic Control Theory
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|}\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right),$
@@ 1.7 Bibliographic Remarks
Control theory is a mature and active discipline with both deep mathematical foundations and a rich history of applications. In terms of application, self-regulating feedback mechanisms that limit the flow of water, much like the modern float valve, were reportedly used in ancient Greece over two thousand years ago. The first mathematical analysis of such systems appeared in a paper by James Clerk Maxwell on the topic of governors [MAX68], shortly after Maxwell had published his treatise on the unification of electricity and magnetism. For an enlightening historical account, see [FZ03].
@@
Online nonstochastic control theory and the Gradient Perturbation Control algorithm were proposed in [ABH+19b]. A flurry of novel methods and regret bounds for online nonstochastic control have appeared in recent years, many of which are surveyed in [HS21]. Notable among them are logarithmic regret algorithms [AHS19c, FS20, SIM20], nonstochastic control for unknown systems [HKS20], black-box online nonstochastic control [CH21], and nonstochastic control with partial observability [SSH20].
@@
In this chapter, we return to the general problem of control from Definition 1.2: choosing a sequence of controls to minimize a sequence of cost functions. Before moving on to algorithms for control, it is useful to ask what kinds of goals are natural and what kinds of guarantees are even possible for a dynamical system.
To that end, we consider several examples of dynamical systems and introduce a few basic solution concepts, including equilibrium, stability, stabilizability, and controllability. These examples also illustrate the kinds of observation models and control objectives that arise in real applications.
@@ 2.1.1 Medical Ventilator
In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m^{3}/s}$) into the lungs of a patient. This inflow approximately determines the volume of the lungs $v_{t}$ via the simple relation
@@
where $\Delta$ is the time elapsed between successive actions of the discrete-time control algorithm. Being difficult to directly measure or estimate, the quantity $v_{t}$ is a hidden or latent variable from the controller’s perspective. Instead, the controller observes the air pressure $p_{t}$. A simplified physical model relating pressure and volume is
$p_{t}=C_{0}+C_{1}v_{t}^{-1/3}+C_{2}v_{t}^{5/3}.$
This relation can be derived from stress–strain considerations by modeling the left and right lungs, connected by the trachea, as in a two-balloon experiment.
The objective of the control algorithm is to follow an ideal pressure trajectory $p_{t}^{\star}$ prescribed by the physician as closely as possible, for example via the stage cost
$c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{\star}|^{2}+|u_{t}|^{2},$
while penalizing large airflow rates so as to limit patient discomfort. See Figure 2.1 for a schematic representation of airflow when a lung is connected to a mechanical ventilator.
@@ 2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving along the real line. The state $\mathbf{x}_{t}\in{\mathbb{R}}^{2}$ is two-dimensional, with the first coordinate indicating position and the second denoting velocity. The control algorithm chooses a scalar actuation $u_{t}\in{\mathbb{R}}$ that directly affects the velocity coordinate at each time step; after an appropriate normalization, $u_{t}$ may be interpreted as a force or acceleration input applied over one sampling interval. The dynamics equation can be written as
@@
The position is linearly related to the velocity through the sampling period $\Delta$. Optionally, the state is also affected by external perturbations $\mathbf{w}_{t}$ outside the influence of the controller.
The goal here is to ensure that the point object arrives and then comes to rest at some coordinate $x^{\star}\in\mathbb{R}$ as quickly as possible, while also not expending too much energy. A natural stage cost is
$c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{\star}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
@@
Figure 2.2: Double integrator illustration showing state coordinate $x_{1}(t)$ (position), state coordinate $x_{2}(t)$ (velocity), and control input $u(t)$.
@@ 2.1.3 Pendulum Swing-Up
The swing-up pendulum is one of the simplest examples of rotational motion. It consists of a point mass $m$ attached to the end of a massless rod of length $l$, with the other end fixed at a pivot. The system is subject to gravity with acceleration $g$, and the controller applies a torque $u_{t}\in\mathbb{R}$ at each time step. If $\theta_{t}$ denotes the angle of the rod measured from the downward vertical, then a discretization of the continuous-time dynamics gives
$\theta_{t+1}$ $=\theta_{t}+\Delta\dot{\theta}_{t}\ \ ,\ \ \ \dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\left(\frac{u_{t}}{ml^{2}}-\frac{g}{l}\sin\theta_{t}\right).$
Thus the state is $\mathbf{x}_{t}=[\theta_{t},\dot{\theta}_{t}]^{\top}$, and the dynamics are nonlinear because of the term $\sin\theta_{t}$.
PivotmmAngle $\theta_{t}$$m\mathbf{g}$$u_{t}$$l$
Figure 2.3: Pendulum swing-up illustration with marked angle $\theta_{t}$, gravitational force mgmg, control torque $u_{t}$, mass $m$, and rod length $l$.
The objective here is to bring the rod to rest in the upright position $\theta=\pi$, using as little torque as possible. A simple stage cost is
@@
When interpreting this objective globally, the angular error should be understood modulo $2\pi$.
2.1.4 Autoregressive Language Models
A modern example of a dynamical system is an autoregressive language model. Let $\mathcal{V}$ be a finite vocabulary, and let $y_{t}\in\mathcal{V}$ denote the token observed at time $t$. The model maintains a hidden state $h_{t}$ that summarizes the preceding context. For a recurrent neural network, one may take $h_{t}\in\mathbb{R}^{d}$ to be a fixed-dimensional latent vector. For a transformer, it is more natural to regard $h_{t}$ as the model’s internal cache or prefix representation associated with the sequence $y_{1:t}$; this state is typically very high-dimensional and is not directly observed.
A simple state-space description is
$h_{t+1}=f_{\theta}(h_{t},e(y_{t})),$
where $e(y_{t})$ is the embedding of token $y_{t}$ and $f_{\theta}$ is a learned nonlinear update map. From the hidden state, the model produces a probability distribution over the next token,
$p_{t+1}=g_{\theta}(h_{t+1})\in\Delta(\mathcal{V}),$
where $\Delta(\mathcal{V})$ denotes the probability simplex over the vocabulary.
The objective in next-token prediction is to assign high probability to the realized next token $y_{t+1}$. A natural stage cost is therefore the log loss
$c_{t}(h_{t},y_{t+1})=-\log p_{t+1}(y_{t+1}).$
Equivalently, one may view training as minimizing the sum of these stage costs over time.
This example illustrates several features that are common in modern machine learning systems. The hidden state is extremely large and latent, the observations are discrete symbols rather than physical measurements, and the dynamics are highly nonlinear. At generation time, the model recursively predicts a distribution over the next token, chooses or samples a token from that distribution, appends it to the context, and repeats the process.
@@ 2.1.5 Epidemiological Models
The COVID-19 pandemic renewed public interest in policy interventions for controlling the spread of infectious diseases. A basic compartmental model is the SIR model. Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the fractions of the population that are susceptible, infected, and removed (recovered or immune), respectively, so that $S_{t}+I_{t}+R_{t}=1$. Let $u_{t}\geq 0$ denote the vaccination effort at time $t$, and let $\alpha\in[0,1]$ denote vaccine efficacy. A discrete-time controlled SIR model is
$S_{t+1}$ $=S_{t}-\beta S_{t}I_{t}-\alpha u_{t},$
$I_{t+1}$ $=I_{t}+\beta S_{t}I_{t}-\gamma I_{t},$
$R_{t+1}$ $=R_{t}+\gamma I_{t}+\alpha u_{t}.$
Here, $\beta$ measures transmissibility, $\gamma$ is the recovery rate, and $\alpha u_{t}$ is the fraction of susceptible individuals effectively moved into the removed/immune compartment by vaccination. Under these dynamics, the identity $S_{t}+I_{t}+R_{t}=1$ is preserved whenever it holds initially.
A natural control objective is to keep infections small while also limiting intervention cost, for example through the stage cost
$c_{t}(S_{t},I_{t},R_{t},u_{t})=I_{t}^{2}+\lambda u_{t}^{2},$
for some tradeoff parameter $\lambda>0$, subject to $u_{t}\geq 0$ and $\alpha u_{t}\leq S_{t}$.
Figure 2.4: Typical trajectory of a continuous-time SIR model without vaccination, with parameters $\beta=0.3$ and $\gamma=0.1$.
These examples span hidden-state and fully observed systems, linear and nonlinear dynamics, and a variety of natural control objectives. We next abstract away from the application details and formalize several general solution concepts.
@@ 2.2 Solution Concepts for Dynamical Systems
Before considering how to control a dynamical system, we should decide what success means. In this section, we discuss several solution concepts, including equilibrium, stability, stabilizability, and controllability.
Given a dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t})$, its deterministic counterpart
$\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t})=f(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{0})$
models the evolution of the system in the absence of perturbations. The autonomous system
$\mathbf{x}_{t+1}=f(\mathbf{x}_{t})=f(\mathbf{x}_{t},\mathbf{0},\mathbf{0})$
describes the state transitions in the absence of both control inputs and perturbations. In both cases, we use the same symbol $f$ when the omitted arguments are understood to be zero.
@@ 2.2.1 Equilibrium of Dynamical Systems
An example of an equilibrium is a ball sitting at rest at the bottom of a pit. If the ball is moved slightly, it returns to the bottom of the pit. Thus this equilibrium is both stable and attractive.
If a ball is at rest on the top of a mountain, then pushing it in any direction causes it to roll away. This is an equilibrium that is unstable. We now formalize these notions.
@@
A point $\mathbf{x}_{e}$ is an equilibrium point of an autonomous dynamical system $\mathbf{x}_{t+1}=f(\mathbf{x}_{t})$ if and only if $f(\mathbf{x}_{e})=\mathbf{x}_{e}$.
The equilibrium $\mathbf{x}_{e}$ is stable if for every $\varepsilon>0$ there exists $\delta>0$ such that
$\|\mathbf{x}_{0}-\mathbf{x}_{e}\|<\delta\quad\Longrightarrow\quad\|\mathbf{x}_{t}-\mathbf{x}_{e}\|<\varepsilon\qquad\text{for all }t\geq 0.$
The equilibrium $\mathbf{x}_{e}$ is attractive if there exists $\delta>0$ such that
$\|\mathbf{x}_{0}-\mathbf{x}_{e}\|<\delta\quad\Longrightarrow\quad\lim_{t\to\infty}\mathbf{x}_{t}=\mathbf{x}_{e}.$
The equilibrium $\mathbf{x}_{e}$ is asymptotically stable if it is both stable and attractive. If the attraction property holds for every initial state in the state space, we say that the equilibrium is globally attractive; if it is also stable, we say that it is globally asymptotically stable.
@@
Clearly $\mathbf{0}$ is an equilibrium point, since $A\mathbf{0}=\mathbf{0}$ for every matrix $A$. If every eigenvalue of $A$ has modulus strictly less than one, then $A^{t}\mathbf{x}_{0}\to\mathbf{0}$ for every initial state $\mathbf{x}_{0}$. Thus, $\mathbf{0}$ is a globally asymptotically stable equilibrium.
In general, determining or even describing equilibrium sets and long-run behavior can be difficult, as the examples later in this chapter illustrate.
@@ 2.2.2 Stabilizability and Controllability
Two fundamental properties of dynamical systems are stabilizability and controllability.
@@
A dynamical system is asymptotically stabilizable if there exists a control policy $\pi$ such that, for every initial state, the deterministic closed-loop trajectory satisfies
$\lim_{t\to\infty}\mathbf{x}_{t}^{\pi}=\mathbf{0}.$
@@
A dynamical system is controllable if, for every initial state $\mathbf{x}_{0}$ and every target state $\mathbf{x}^{\star}$, there exist a finite horizon $T$ and controls $\mathbf{u}_{0},\ldots,\mathbf{u}_{T-1}$ such that the deterministic counterpart of the system satisfies $\mathbf{x}_{T}=\mathbf{x}^{\star}$.
Controllability implies finite-time reachability of the zero state. If the zero state is an equilibrium that can be maintained once reached, then controllability also implies asymptotic stabilizability.
Some texts reserve the word stabilizable for a weaker algebraic property of linear systems. Here we use the more direct trajectory-based notion above, since it is convenient for our purposes.
In the next subsection, we introduce a more quantitative robust-boundedness notion that is useful in nonstochastic control.